# Examining Racial Stereotypes in YouTube Autocomplete Suggestions

### v2

@@ Abstract
Autocomplete is a popular search feature that predicts queries based on user input and guides users to a set of potentially relevant suggestions. In this study, we examine how YouTube autocompletes serve as an information source for users exploring information about race. We perform an algorithm output audit of autocomplete suggestions for input queries about four racial groups and examine the stereotypes they embody. Using critical discourse analysis, we identify five major sociocultural contexts in which racial biases manifest—Appearance, Ability, Culture, Social Equity, and Manner. Our results show evidence of aggregated discrimination and interracial tensions in the autocompletes we collected and highlight their potential risks in othering racial minorities. We call for urgent innovations in content moderation policy design and enforcement to address these biases in search outputs.
@@
**footnotetext: These authors contributed equally to this work
[…]
Shagun Jhaver, Rutgers University School of Communication & Information, New Brunswick, NJ
@@ 1 Introduction
Most prior research on autocomplete suggestions centers on algorithmic outputs from general-purpose search engines, especially Google Search (Al-Abbas et al., 2020; Baker and Potts, 2013; Leidinger and Rogers, 2023; Roy and Ayalon, 2020). This research focus seems reasonable because Google Search has been the largest global search engine since the early 2000’s. [^2] https://www.britannica.com/topic/Google-Inc However, the use of autocomplete to assist query inputs goes far beyond Google searching. For instance, YouTube, the platform we focus on in this study, has incorporated the autocomplete feature in its search function for many years. Indeed, if we count the number of searches made across all platforms, YouTube emerges as the second-largest search engine after Google (Smith, 2020), offering all its results in a video format. However, YouTube autocomplete has attracted little scholarly attention, perhaps because of YouTube’s public perception as a video-sharing platform rather than a search engine.
Unlike search engines, YouTube allows users to not just view, but also create content on its site (Wotanis and McMillan, 2014). Prior research on YouTube’s participatory culture has characterized it as a critical form of expression, “self-promotion, escapism, and utter play.” (Boxman-Shabtai, 2019) However, it is possible that YouTube’s complex cultural fabric comprising both professional (i.e., traditional media) and amateur (i.e., independent individuals) content production (Boxman-Shabtai, 2019) induces the types of biases that do not appear on search engines.
@@
Specifically, we explore what information YouTube users in the United States (U.S.) encounter through the search autocomplete suggestions. Since misinformation or limited information about racial groups, along with media stereotypes, can shape individuals’ perceptions of race (Covington, 2010), we pose the following research question: what information does YouTube provide to users exploring information about racial topics, and how may that information shape users’ understandings of race? We focus on race here to extend prior conversations on the media’s construction of race (Chen, 1996; Squires, 2014).
Importantly, the autocomplete feature not only provides information but may also contribute to the propagation of racial stereotypes. We perform an algorithmic output audit of YouTube autocomplete suggestions regarding four racial groups—Whites, Blacks, Asians, and Hispanics—to surface such stereotypes. We contend that discriminatory race-related outputs generated by algorithms, including autocomplete suggestions, should be interrogated as the most visible signs of deeper issues related to racial biases inherent in sociotechnical systems driving such algorithms.
@@ 2.1 Racial Stereotypes
Historically, Blacks or African Americans have been negatively stereotyped with respect to a variety of sociocultural characteristics, including their personality, hygiene and criminal conduct. Messages from mass media and political communications often reinforce these stereotypes (Hurwitz and Peffley, 1997). Some scholars maintain that these distorted perceptions of Blacks are rooted in slavery (Lintner, 2004) and Christian mythology (Miller, 1995). Whatever their origin, such racial stereotypes are dangerous because empowered non-minorities can exploit them to justify their privilege and continue to perpetuate harm and discrimination against ‘others.’ For instance, racial segregation in the U.S. was historically based on biased perceptions about Blacks’ personalities, intelligence, and appearance (Lintner, 2004).
@@ 2.2 Algorithmic Bias, Search Critiques, and Algorithm Audits
The research presented here can be placed within the broader critique of algorithmic biases in content moderation, though we focus on search feature outputs rather than AI-assisted regulation of potentially inappropriate social media posts (Jhaver et al., 2019). Specifically, this article aligns with approaches that aim to assess the extent to which search outputs are: problematic; characterized by offensive, stereotypical, inappropriate, and discriminatory terms; and/or racist (Rogers, 2023). Our method of probing YouTube’s search feature with specific queries could be positioned alongside approaches like algorithm audits (Sandvig et al., 2014) and ethical hacking practices that search for vulnerabilities (Roberts, 2019).
@@
Our work contributes to this line of research by deploying a novel methodological approach and conceptual focus to understanding the prevalence of racial biases in YouTube search autocompletes. While prior studies have primarily focused on only the text of autocomplete suggestions (Al-Abbas et al., 2020; Baker and Potts, 2013; Leidinger and Rogers, 2023; Roy and Ayalon, 2020), we analyze autocomplete texts in conjunction with the video results for each suggestion. This helps us present a more nuanced understanding of how YouTube users’ interactions with its search feature may influence their perspectives. Our work is the first to deploy critical discourse analysis (CDA) in addition to inductive analysis to uncover not only how language and discourse practices of autocompletes reflect problematic socio-cultural norms but also how they reinforce hegemonic whiteness and racist ideologies.
@@ 3.2 Data Collection
Analyzing the data samples from three different waves and five different states, we found minimal differences among them. We integrated these samples and removed duplicate results to create a combined dataset for further inquiry. While this step precluded prioritizing more geographically and temporally stable autocompletes, it allowed us to work with a more comprehensive set of suggestions. Examining the resulting dataset, we also noticed that some autocomplete suggestions were not related to the racial groups. Thus, we further removed the irrelevant autocomplete suggestions based on whether the autocomplete refers to non-human rather than non-human subjects. After removing duplicates, our dataset consisted of a total of 241 unique autocompletes.
@@ 3.3 Data Analysis
We began by developing a situated understanding of each autocomplete suggestion by examining about ten videos displayed as top search results when selecting that suggestion. This examination helped us recognize the contextual and cultural meaning of autocomplete suggestions and we leveraged this nuanced understanding to develop our open codes. This careful, context-driven coding process was important because some of the autocomplete results were closely related to memes or video content in their corresponding search results. For instance, the autocomplete “when an asian is kidnapped” was related to comedy content created by an Asian YouTuber. Thus, it was important to search for and understand the situated meanings of autocompletes rather than interpret them in isolation.
@@
Then, we developed the common themes arising from our open codes and their corresponding autocomplete data by grouping similar codes into conceptual categories. At this point in the process, when any pre-existing categories seemed not to directly align with an emergent theme, we carefully considered whether to introduce a new category, revise an existing category, and/or create a subcategory. This process was iterative and we compared our emerging categories with one another and with the underlying data at each stage of our analysis. Some autocompletes represented miscellaneous sociocultural contexts that did not warrant inclusion in or creation of a separate category; we excluded such autocompletes from further analysis. In total, we grouped 217 autocompletes into a two-tier category structure.
@@
Throughout the process of qualitative coding and categorization of autocomplete results, we incorporated critical discourse analysis (CDA). CDA is a systematic approach to examining a text, event, or discursive practice to identify how it originates from and is affected by power relations (Fairclough, 1993). We employed this framework to analyze each autocomplete as concurrently a text (language produced in a discursive event), a discourse practice (the production, distribution, and consumption of a text), and a sociocultural practice (hegemonic struggles that ideologically shape a discursive practice) (Fairclough, 1995). This meant that during data analysis, our focus was less on examining the lexical meanings of each autocomplete suggestion than on the hegemonic function the suggestion performed or indicated within the sociocultural practices represented by our five main categories. Using this approach, we aimed to identify various facets of each discursive event, thereby revealing particular power dynamics reflected in that event. We engaged in regular discussions about how the autocomplete suggestions in our data reproduce and legitimize social inequalities, and our coding sought to identify various forms of domination and biases.
@@ 4.1 Appearance
Personal Hygiene. This category is notable because only Blacks were subject to autocompletes related to personal hygiene. These autocompletes referenced Black people washing their hair or beard (“how often should a black man wash his hair,” “how often should a black woman wash her hair,” “how often should a black man wash his beard”) and shaving (“should a black man shave everyday”). Figure 3 shows the preview of a video resulting from the suggested autocompletes, which contains advice for preventing white flakes in hair. [^3] https://www.youtube.com/watch?v=CZCuI4x1sUc Another autocomplete asked, “do black people get lice.” It is possible that these autocompletes reflect highly sought-after searches by Black people for the information needed to take care of textured hair. However, these questions are raised only for the Black group, raising the possibility that they reflect the racial stereotypes of Black people having poor hygiene. Historically, a deep-rooted racial stereotype maintained that dark or non-white skin was associated with uncleanness (Brown, 2009); indeed, this stereotype served as a rationale for racial segregation in the U.S. (Lintner, 2004). These YouTube autocompletes suggest that the negative stereotypes about Blacks’ hygiene unfortunately persist.
@@
Skin Tone. Similar to ‘Personal Hygiene,’ Blacks were the main subjects of autocompletes in the ‘Skin Tone’ category. Autocompletes related to skin tones included “can a black person become fair.” These bleaching-related autocompletes reflect both an interest in lighter skin and the implicit negative view of darker skin (Hunter, 2011). We observed no autocompletes related to the action of changing natural skin tones for other races. Other autocompletes show how Black skin is perceived as unique and distinct from other races. Examples include autocompletes asking about the need for Black people’s skin protection (“should black people wear sunscreen,” “do black people get sunburn”) and changes in their natural skin tones (“can black people blush,” “can black people tan”). These results suggest both an otherization and a general lack of awareness about Black skin characteristics.
@@ 4.2 Ability
Financial. Only Blacks were subjects of autocompletes that belonged to financial ability. All autocompletes here were associated with a lack of financial resources and acumen, which reinforces poor Blacks as the representative public image of poverty. Interestingly, this is consistent with how network TV news and weekly news magazines have historically grossly overrepresented Blacks in their portrayal of poor people as a whole (Gilens, 1996). Autocompletes in this category promoted the stereotype that Blacks cannot purchase or own expensive products, such as fancy cars or clothes (“black man cannot own g wagon,” “black woman can’t buy a dress,” “a black man can’t have a suitcase”).
@@ 4.3 Culture
Second, autocompletes associated with culturally specific attire and beauty standards also indicated instances of cultural appropriation, where members of the majority group coopted traditional garments, hairstyles, or makeup practices of other minority groups. For example, autocompletes related to garments or hairstyles for White groups included traditional styles with deep origins in Asian or Black culture, e.g., “can a white person wear a kimono,” “can a white person wear a durag,” “can white girls get braids,” and “can a white boy get waves” (Figure 5). [^5] https://www.youtube.com/watch?v=gSZ8U64EL-Q Additionally, we found instances of blackfishing, i.e., the practices of modifying one’s appearances physically or digitally or using terms and language patterns linked to Ebonics for racial commodification (Stevens, 2021, p.1). For example, we found autocompletes for White groups that included references to the makeup style of White people attempting to look like Blacks (“the white woman who turned black”).
@@ 4.4 Social Equity
Diversity/Inclusion. Whites, Blacks, and Hispanics had autocompletes associated with the issues of diversity and inclusion. Curiously, most of these autocompletes referred to the underrepresentation of racial minorities in popular culture (“can a black man be captain america,” “is there a hispanic disney princess” Figure 7). [^8] https://www.youtube.com/watch?v=AQLUeshn3nE This reflects a cultural appetite for seeing racial diversity and encouraging inclusive casting decisions in mainstream media. While whitewashing has historically been salient in the entertainment industry, the percentage of White characters depicted on-screen has slightly declined in recent years, while the depiction of Asian characters has significantly increased (Smith et al., 2023). However, no notable changes have occurred in the prevalence of Black or Hispanic characters over the same period (Smith et al., 2023). Seemingly in response to these trends, YouTube autocompletes in this category reveal frustrations about persistent inequalities in the racial representation of Black and Hispanic character portrayals.

### v3

@@ Abstract
Autocomplete is a popular search feature that predicts queries based on user input and guides users to a set of potentially relevant suggestions. In this study, we examine what YouTube autocompletes suggest to users seeking information about race on the platform. Specifically, we perform an algorithm output audit of autocomplete suggestions for input queries about four racial groups and examine the stereotypes they embody. Using critical discourse analysis, we identify five major sociocultural contexts in which racial information appears—Appearance, Ability, Culture, Social Equity, and Manner. We found that the participatory nature of YouTube produces a multifaceted representation of race-related content in its search outputs, characterized by enduring historical biases, aggregated discrimination, and interracial tensions, while simultaneously depicting minority resistance and aspirations of a post-racial society. We call for innovations in content moderation policy design and enforcement to address existing racial harms in YouTube search outputs.
@@ 1 Introduction
Most prior research on autocomplete suggestions centers on algorithmic outputs from general-purpose search engines, especially Google Search (Al-Abbas et al., 2020; Baker and Potts, 2013; Leidinger and Rogers, 2023; Roy and Ayalon, 2020). This research focus seems reasonable because Google Search has been the largest global search engine since the early 2000’s. [^2] https://www.britannica.com/topic/Google-Inc However, the use of autocomplete to assist query inputs goes far beyond Google searching. For instance, YouTube, the platform we focus on in this study, has incorporated the autocomplete feature in its search function for many years. Indeed, if we count the number of searches made across all platforms, YouTube emerges as the second-largest search engine after Google (Smith, 2020), offering all its results in a video format. While Google does not publicly describe the extent to which it uses the same data and algorithms to drive search autocomplete suggestions on Google and YouTube, official documentation suggests that YouTube autocomplete predictions are at least partly shaped by content specific to YouTube, including users’ video watching history (YouTube, 2025). However, YouTube autocomplete has attracted little scholarly attention, perhaps because of YouTube’s public perception as a video-sharing platform rather than a search engine.
Unlike search engines, YouTube allows users to not just view, but also create content on its site (Wotanis and McMillan, 2014). Prior research on YouTube’s participatory culture has characterized it as a critical form of expression, “self-promotion, escapism, and utter play” (Boxman-Shabtai, 2019). However, it is possible that YouTube’s complex cultural fabric comprising both professional (i.e., traditional media) and amateur (i.e., independent individuals) content production (Boxman-Shabtai, 2019) induces the types of biases that do not appear on search engines. At the same time, independent YouTubers have the potential to create and share content that counteracts traditional biases and influence search outputs in positive ways.
@@
Specifically, we explore what information YouTube users in the United States (U.S.) encounter through the search autocomplete suggestions. Since misinformation or limited information about racial groups, along with media stereotypes, can shape individuals’ perceptions of race (Covington, 2010), we pose the following research question: what suggestions does the YouTube search feature provide to users exploring information about racial topics, and how may such suggestions shape users’ understandings of race? We focus on race here to extend prior conversations on the media’s construction of race (Chen, 1996; Squires, 2014; Walters et al., 2024).
Importantly, the autocomplete feature not only provides information but may also contribute to the propagation of racial stereotypes. We perform an algorithmic output audit of YouTube autocomplete suggestions regarding four racial groups—Whites, Blacks, Asians, and Hispanics—to surface such stereotypes. At the same time, we also attend to positive representations of racial minorities and attempts to counteract racial stigmas in these autocomplete suggestions.
@@ 2.1 Racial Stereotypes
Historically, Blacks or African Americans have been negatively stereotyped with respect to a variety of sociocultural characteristics, including their personality, hygiene, and criminal conduct. Messages from mass media and political communications often reinforce these stereotypes (Hurwitz and Peffley, 1997). Some scholars maintain that these distorted perceptions of Blacks are rooted in slavery (Lintner, 2004) and Christian mythology (Miller, 1995). Whatever their origin, such racial stereotypes are dangerous because empowered non-minorities can exploit them to justify their privilege and continue to perpetuate harm and discrimination against ‘others.’ For instance, racial segregation in the U.S. was historically based on biased perceptions about Blacks’ personalities, intelligence, and appearance (Lintner, 2004).
@@ 2.2 Algorithmic Bias, Search Critiques, and Algorithm Audits
The research presented here can be placed within the broader critique of algorithmic biases in content moderation, though we focus on search feature outputs rather than AI-assisted regulation of potentially inappropriate social media posts (Jhaver et al., 2019). Specifically, this article aligns with approaches that aim to assess the extent to which search outputs are: problematic; characterized by offensive, stereotypical, inappropriate, and discriminatory terms; and/or racist (Rogers, 2023). Unlike these approaches, we also document how the search outputs serve to dismantle racist stereotypes. Our method of probing YouTube’s search feature with specific queries could be positioned alongside approaches like algorithm audits (Sandvig et al., 2014) and ethical hacking practices that search for vulnerabilities (Roberts, 2019).
@@
Our work contributes to this line of research by examining the prevalence of racial biases in YouTube search autocompletes. While prior studies have primarily focused on only the text of autocomplete suggestions (Al-Abbas et al., 2020; Baker and Potts, 2013; Leidinger and Rogers, 2023; Roy and Ayalon, 2020), we analyze autocomplete texts in conjunction with the video results for each suggestion. This helps us present a more nuanced understanding of how YouTube users’ interactions with its search feature may influence their perspectives. Our work is the first to deploy critical discourse analysis (CDA) in addition to inductive analysis to uncover not only how language and discourse practices of autocompletes reflect problematic socio-cultural norms but also how they reinforce hegemonic whiteness and racist ideologies.
@@ 3.2 Data Collection
Analyzing the data samples from three different waves and five different states, we found minimal differences among them. We integrated these samples and removed duplicate results to create a combined dataset for further inquiry. While this step precluded prioritizing more geographically and temporally stable autocompletes, it allowed us to work with a more comprehensive set of suggestions. Examining the resulting dataset, we also noticed that some autocomplete suggestions were not related to the racial groups. Thus, we removed the irrelevant autocomplete suggestions based on whether the autocomplete refers to non-human rather than human subjects. After removing duplicates, our dataset consisted of a total of 241 unique autocompletes.
@@ 3.3 Data Analysis
We began by developing a situated understanding of each autocomplete suggestion by examining about ten videos displayed as top search results when selecting that suggestion. This examination helped us recognize the contextual and cultural meaning of autocomplete suggestions, and we leveraged this nuanced understanding to develop our open codes. Since we were primarily interested in the autocomplete texts, we did not create open codes to describe the video content; instead, we created codes and memos that delineated the racial biases, tensions, and struggles suggested by the corresponding autocompletes and the sociocultural contexts in which they occurred. This careful, context-driven coding process was important because some of the autocomplete results were closely related to memes or video content in their corresponding search results. For instance, the autocomplete “when an asian is kidnapped” was related to comedy content created by an Asian YouTuber. Thus, it was important to search for and understand the situated meanings of autocompletes rather than interpret them in isolation. However, in many cases, the resulting videos did not directly relate to or provide any additional context for the autocomplete text; we suspect that in such cases, the autocomplete texts could have been influenced by the search patterns of YouTube users rather than the content of YouTube videos.
@@
Then, we developed the common themes arising from our open codes and their corresponding autocomplete data by grouping similar codes into conceptual categories. At this point in the process, when any pre-existing categories seemed not to directly align with an emergent theme, we carefully considered whether to introduce a new category, revise an existing category, and/or create a subcategory. This process was iterative and we compared our emerging categories with one another and with the underlying data at each stage of our analysis. Some autocompletes represented miscellaneous contexts that did not warrant inclusion in or creation of a separate category; we excluded such autocompletes from further analysis. In total, we grouped 217 autocompletes into a two-tier category structure, where the main categories represented the overarching sociocultural contexts in which racial biases manifest, and the subcategories represented more specific discursive events within these contexts. The five main categories that emerged through this process included Appearance, Ability, Culture, Social Equity, and Manner (see Table 3).
@@
Throughout the process of qualitative coding and categorization of autocomplete results, we incorporated critical discourse analysis (CDA). CDA is a systematic approach to examining a text, event, or discursive practice to identify how it originates from and is affected by power relations (Fairclough, 1993). We employed this framework to analyze each autocomplete as concurrently a text (language produced in a discursive event), a discourse practice (the production, distribution, and consumption of a text), and a sociocultural practice (hegemonic struggles that ideologically shape a discursive practice) (Fairclough, 1995). This meant that during data analysis, our focus was less on examining the lexical meanings of each autocomplete suggestion than on the hegemonic function the suggestion performed or indicated within the sociocultural practices represented by our five main categories. For many autocompletes, our attention to their corresponding video results helped us characterize their text production, consumption, and underlying hegemonic struggles. Using this approach, we aimed to identify various facets of each discursive event, thereby revealing particular power dynamics reflected in that event. We engaged in regular discussions about how the autocomplete suggestions in our data reproduce and legitimize social inequalities, and our coding sought to identify various forms of domination, biases, and resistance.
@@ 4.1 Appearance
Personal Hygiene. This category is notable because only Blacks were subject to autocompletes related to personal hygiene. These autocompletes referenced Black people washing their hair or beard (“how often should a black man wash his hair,” “how often should a black woman wash her hair,” “how often should a black man wash his beard”) and shaving (“should a black man shave everyday”). The resulting videos featured creators presenting their shower or shaving routines, e.g., Figure 3 shows the preview of a video resulting from the suggested autocompletes, which contains advice for preventing white flakes in hair. [^3] https://www.youtube.com/watch?v=CZCuI4x1sUc Another autocomplete asked, “do black people get lice.” It is possible that these autocompletes reflect highly sought-after searches by Black people for the information needed to take care of textured hair. However, these questions are raised only for the Black group, raising the possibility that they reflect the racial stereotypes of Black people having poor hygiene. Historically, a deep-rooted racial stereotype maintained that dark or non-white skin was associated with uncleanness (Brown, 2009); indeed, this stereotype served as a rationale for racial segregation in the U.S. (Lintner, 2004). These YouTube autocompletes suggest that the negative stereotypes about Blacks’ hygiene still persist.
@@
Skin Tone. Similar to ‘Personal Hygiene,’ Blacks were the main subjects of autocompletes in the ‘Skin Tone’ category. Autocompletes related to skin tones included “can a black person become fair.” These bleaching-related autocompletes reflect both an interest in lighter skin and the implicit negative view of darker skin (Hunter, 2011). We observed no autocompletes related to the action of changing natural skin tones for other races. Other autocompletes show how Black skin is perceived as unique and distinct from other races. Examples include autocompletes asking about the need for Black people’s skin protection (“should black people wear sunscreen,” “do black people get sunburn”) and changes in their natural skin tones (“can black people blush,” “can black people tan”). These results suggest both an otherization and a general lack of awareness about Black skin characteristics. However, some videos for these autocompletes showed creators dismantling stereotypes about Black skin, e.g, in one such video, [^4] https://www.youtube.com/shorts/tMkpBnmNy7o a Black creator applied blush to her skin.
@@ 4.2 Ability
Financial. Only Blacks were subjects of autocompletes that belonged to financial ability. All autocompletes here were associated with a lack of financial resources and acumen, which reinforces poor Blacks as the representative public image of poverty. Interestingly, this is consistent with how network TV news and weekly news magazines have historically grossly overrepresented Blacks in their portrayal of poor people as a whole (Gilens, 1996). Autocompletes in this category reflected the stereotype that Blacks cannot purchase or own expensive products, such as fancy cars or clothes (“black man cannot own g wagon,” “black woman can’t buy a dress,” “a black man can’t have a suitcase”).
@@ 4.3 Culture
Second, autocompletes associated with culturally specific attire and beauty standards also indicated instances of cultural appropriation, where members of the majority group coopted traditional garments, hairstyles, or makeup practices of other minority groups. For example, autocompletes related to garments or hairstyles for White groups included traditional styles with deep origins in Asian or Black culture, e.g., “can a white person wear a kimono,” “can a white person wear a durag,” “can white girls get braids,” and “can a white boy get waves” (Figure 5). [^6] https://www.youtube.com/watch?v=gSZ8U64EL-Q Additionally, we found instances of blackfishing, i.e., the practices of modifying one’s appearances physically or digitally or using terms and language patterns linked to Ebonics for racial commodification (Stevens, 2021, p.1). For example, we found autocompletes for White groups that included references to White people attempting to look like Blacks (e.g., “the white woman who turned black”). Some videos linked by these autocompletes displayed the television coverage of Martina Big, a German model who underwent a perma-tanning procedure to darken her skin color.
@@ 4.4 Social Equity
Diversity/Inclusion. Whites, Blacks, and Hispanics had autocompletes associated with the issues of diversity and inclusion. Curiously, most of these autocompletes referred to the underrepresentation of racial minorities in popular culture (“can a black man be captain america,” “is there a hispanic disney princess”; see Figure 7). [^9] https://www.youtube.com/watch?v=AQLUeshn3nE This reflects a cultural appetite for seeing racial diversity and encouraging inclusive casting decisions in mainstream media. While whitewashing has historically been salient in the entertainment industry, the percentage of White characters depicted on-screen has slightly declined in recent years, while the depiction of Asian characters has significantly increased (Smith et al., 2023). However, no notable changes have occurred in the prevalence of Black or Hispanic characters over the same period (Smith et al., 2023). Seemingly in response to these trends, YouTube autocompletes in this category reveal frustrations about persistent inequalities in the racial representation of Black and Hispanic character portrayals.